OpenStudy (anonymous):

The solution set of 2x^2=50 is?

jimthompson5910 (jim_thompson5910):

2x^2=50 x^2=50/2 x^2 = 25 x = +-sqrt(25) x = 5 or x = -5

OpenStudy (anonymous):

x=\[x =\sqrt{25} =5\]

OpenStudy (anonymous):

yeah positive and negative 5